A particle of mass 20 g is thrown vertically upwards with a speed of 10 m/s. Find the work done by the force of gravity during the time the particle goes up.

Grade:12th pass

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
Suppose the particle reaches a maximum height h. As the velocity at the highest point is zero, we have

0 = u^{2} - 2gh

or,

h = \frac{u^{2}}{2g}

The work done by the force of gravity is

-mgh = -mg\frac{u^{2}}{2g} = -\frac{1}{2}mu^{2}

= -\frac{1}{2}(0.02 kg )\times (10 m/s^{2}) = -1.0 J